# Tallahassee, FL 32306-4350, USA

### v1

@@ A.1 $U(3)^{5}$ flavour symmetric SMEFT
$C_{qe}^{[prst]}=C_{qe}\delta_{pr}\delta_{st}\,,\qquad C_{qu}^{(1),(8)[prst]}=C_{qu}^{(1),(8)}\delta_{pr}\delta_{st}\,,\qquad C_{qd}^{(1),(8)[prst]}=C_{qd}^{(1),(8)}\delta_{pr}\delta_{st}\,.$
@@ Appendix B Tables of results
In this appendix we collect a series of tables with the numerical results presented in the figures discussed in the main text, as well as some supplementary results. We report both the $68\%$ HPDI interval for the Wilson coefficients at the different scales considered in the study, as well as the bound on $\Lambda/\sqrt{|C|}$ in units of TeV obtained from the maximum of the $95\%$ HPDI interval for $|C|$. As indicated in the main text, in some cases the $68\%$ or $95\%$ HPDI (from which the bound on $\Lambda/\sqrt{|C|}$ is obtained) touches the edges of the prior dictated by perturbativity. These cases are indicated in the tables in red colour. On the other hand, in those cases where the posterior distribution of a given coefficient does not allow to set a bound, we indicate that using a dash.
@@ B.1 Results for the $U(3)^{5}$ flavour symmetric SMEFT
Table 7: Results of the individual fits for the $U(3)^{5}$ flavour symmetric SMEFT. For each coefficient, we report the $68\%$ HPDI interval, the bound on $\Lambda/\sqrt{|C|}$ in units of TeV obtained taking the maximum of the $95\%$ HPDI interval for $|C|$, and the set of observables that give the strongest constraint (EW stands for Electroweak, F for Flavour, H for Higgs, T for top, and DY for Drell-Yan). Results are shown with RG evolution for $\Lambda=10$ TeV, for $\Lambda=3$ TeV and without RG evolution. A dash denotes that no bound can be set from that particular fit. The meaning of the colour code is discussed in the text.
@@ 10 TeV 3 TeV 1 TeV - noRGE
$C$ $\Lambda/\sqrt{|C|}$ Obs $C$ $\Lambda/\sqrt{|C|}$ Obs $C$ $\Lambda/\sqrt{|C|}$ Obs
$C_{\phi\Box}$ H $-0.8\pm 3.0$ $1.2$ H $-0.03\pm 0.28$ $1.3$ H
$C_{\phi D}$ $0.4\pm 1.0$ $6.4$ EW $0.031\pm 0.088$ $6.5$ EW $0.0022\pm 0.0072$ $7.7$ EW
$C_{\phi G}$ $-0.114\pm 0.082$ $19.4$ H $-0.0103\pm 0.0084$ $18.2$ H $-0.0016\pm 0.0012$ $15.7$ H
$C_{\phi B}$ $-0.07\pm 0.17$ $15.8$ H $-0.009\pm 0.014$ $15.8$ H $-0.0007\pm 0.0014$ $16.8$ H
$C_{\phi W}$ $-0.31\pm 0.51$ $8.6$ H $-0.026\pm 0.047$ $8.7$ H $-0.0031\pm 0.0050$ $8.9$ H
$C_{\phi WB}$ $0.19\pm 0.29$ $11.3$ H $0.018\pm 0.025$ $11.5$ H $0.0015\pm 0.0025$ $12.4$ H
$C_{G}$ $8.5\pm 4.5$ H $1.07\pm 0.63$ $2.0$ H $0.07\pm 0.15$ $1.7$ T
$C_{W}$ $-2.3\pm 3.8$ $3.2$ H $-0.24\pm 0.43$ $2.9$ H $0.05\pm 0.69$ $0.8$ EW
$C_{\phi l}^{(1)}$ $-0.33\pm 0.45$ $9.1$ EW $-0.030\pm 0.039$ $9.2$ EW $-0.0029\pm 0.0039$ $9.8$ EW
$C_{\phi l}^{(3)}$ $-0.67\pm 0.64$ $7.3$ EW $-0.058\pm 0.056$ $7.3$ EW $-0.0061\pm 0.0058$ $7.6$ EW
@@
$C_{\phi d}$ ${\color[rgb]{1,0,0}-10.3\pm 4.4}$ EW $-0.83\pm 0.51$ $2.2$ EW $-0.087\pm 0.051$ $2.3$ EW
$C_{ll}^{[aabb]}$ $-1.7\pm 2.1$ $4.2$ EW $-0.15\pm 0.19$ $4.1$ EW $-0.019\pm 0.021$ $4.0$ EW
$C_{ll}^{[abba]}$ $-0.01\pm 0.96$ $7.3$ EW $-0.005\pm 0.091$ $6.9$ EW $-0.004\pm 0.010$ $6.3$ EW
$C_{qq}^{(1)[aabb]}$ $-5.8\pm 4.6$ $2.7$ EW $-0.62\pm 0.54$ $2.3$ EW $-1.00\pm 0.55$ $0.7$ T
@@
$C$ $\Lambda/\sqrt{|C|}$ $C$ $\Lambda/\sqrt{|C|}$ $C$ $\Lambda/\sqrt{|C|}$ $C$ $\Lambda/\sqrt{|C|}$
$C_{\phi\Box}$ $-4.9\pm 3.2$ $0.9$ $5.8\pm 5.4$
$C_{\phi D}$ $0.036\pm 0.087$ $6.6$ $-0.2\pm 4.1$ $1.0$ $0.4\pm 1.0$ $6.4$
$C_{\phi G}$ $-0.014\pm 0.010$ $16.1$ $0.008\pm 0.028$ $11.9$ $-0.114\pm 0.082$ $19.4$ $-0.04\pm 0.17$ $15.7$
$C_{\phi B}$ $-0.017\pm 0.015$ $13.8$ $-0.17\pm 0.51$ $2.9$ $-0.07\pm 0.17$ $15.8$ $-0.8\pm 2.5$ $4.6$
$C_{\phi W}$ $-0.063\pm 0.046$ $7.6$ $-1.8\pm 1.5$ $1.4$ $-0.31\pm 0.51$ $8.6$ ${\color[rgb]{1,0,0}-11.3\pm 3.8}$
$C_{\phi WB}$ $0.036\pm 0.026$ $10.2$ $-1.3\pm 1.5$ $1.5$ $0.19\pm 0.29$ $11.3$ $-7.15\pm 4.65$
$C_{G}$ $1.21\pm 0.78$ $1.8$ $2.6\pm 2.0$ $1.2$ $8.5\pm 4.5$ ${\color[rgb]{1,0,0}8.4\pm 6.65}$
$C_{W}$ $-0.60\pm 0.43$ $2.5$ $0.5\pm 7.2$ $-2.3\pm 3.8$ $3.2$ $-6.8\pm 8.2$
$C_{\phi l}^{(1)}$ $-0.030\pm 0.039$ $9.2$ $-0.3\pm 1.2$ $1.8$ $-0.33\pm 0.45$ $9.1$ $3.4\pm 3.2$ $3.3$
$C_{\phi l}^{(3)}$ $-0.058\pm 0.056$ $7.3$ $0.78\pm 0.36$ $2.5$ $-0.67\pm 0.64$ $7.3$ $4.91\pm 3.25$ $2.9$
$C_{\phi e}$ $0.052\pm 0.051$ $7.7$ $1.0\pm 2.0$ $1.3$ $0.62\pm 0.59$ $7.6$ ${\color[rgb]{1,0,0}9.8\pm 5.1}$
$C_{\phi q}^{(1)}$ $0.10\pm 0.13$ $5.0$ $-3.2\pm 1.4$ $1.2$ $0.8\pm 1.4$ $5.4$ $-2.3\pm 4.6$ $2.9$
$C_{\phi q}^{(3)}$ $0.086\pm 0.052$ $6.8$ $0.58\pm 0.23$ $3.0$ $1.01\pm 0.58$ $6.7$ $6.285\pm 2.205$ $3.0$
$C_{\phi u}$ $0.13\pm 0.18$ $4.3$ $0.9\pm 2.1$ $1.4$ $1.3\pm 1.7$ $4.7$ ${\color[rgb]{1,0,0}-6.8\pm 7.2}$
$C_{\phi d}$ $-0.83\pm 0.51$ $2.2$ $-4.0\pm 2.8$ $1.0$ ${\color[rgb]{1,0,0}-10.3\pm 4.4}$ ${\color[rgb]{1,0,0}-8.2\pm 6.8}$
$C_{ll}^{[aabb]}$ $-0.15\pm 0.19$ $4.1$ $-1.7\pm 2.1$ $4.2$
$C_{ll}^{[abba]}$ $-0.005\pm 0.091$ $6.9$ $0.58\pm 0.88$ $2.1$ $-0.01\pm 0.96$ $7.3$ $2.0\pm 3.4$ $3.4$
$C_{qq}^{(1)[aabb]}$ $-0.62\pm 0.54$ $2.3$ $-11.1\pm 3.9$ $-5.8\pm 4.6$ $2.7$ ${\color[rgb]{1,0,0}-8.6\pm 6.5}$
$C_{qq}^{(1)[abba]}$ $0.09\pm 0.23$ $4.0$ $-0.3\pm 1.4$ $1.7$ $0.3\pm 2.2$ $4.7$ $3.42\pm 5.12$ $2.7$
$C_{qq}^{(3)[aabb]}$ $0.034\pm 0.050$ $8.3$ $-0.42\pm 0.35$ $2.9$ $0.21\pm 0.40$ $9.9$ $-0.5\pm 2.9$ $4.2$
$C_{qq}^{(3)[abba]}$ $0.25\pm 0.16$ $4.0$ $-0.34\pm 0.71$ $2.3$ $2.0\pm 1.4$ $4.5$ $4.3226\pm 2.9725$ $3.2$
$C_{lq}^{(1)}$ $0.055\pm 0.054$ $7.4$ $-2.9\pm 2.3$ $1.2$ $0.23\pm 0.51$ $8.8$ $-3.58\pm 5.28$ $2.7$
$C_{lq}^{(3)}$ $-0.0035\pm 0.0055$ $24.9$ $-0.13\pm 0.11$ $5.1$ $-0.028\pm 0.058$ $26.7$ $-0.03\pm 0.45$ $10.2$
$C_{ee}$ $-0.10\pm 0.12$ $5.3$ $-2.3\pm 7.5$ $0.9$ $-1.0\pm 1.3$ $5.3$ $-3.4\pm 6.6$
$C_{uu}^{[aabb]}$ $0.47\pm 0.56$ $2.4$ $-9.5\pm 5.6$ $2.2\pm 4.4$ $3.1$
$C_{uu}^{[abba]}$ $0.51\pm 0.40$ $2.6$ $-2.5\pm 2.2$ $1.2$ $4.2\pm 4.1$ $2.8$ ${\color[rgb]{1,0,0}-8.9\pm 6.1}$
@@
$C_{dd}^{[abba]}$ $5.0\pm 5.0$
$C_{eu}$ $0.036\pm 0.038$ $9.0$ $0.9\pm 3.5$ $1.1$ $0.17\pm 0.41$ $10.2$ $0.2\pm 7.5$
$C_{ed}$ $-0.10\pm 0.12$ $5.2$ ${\color[rgb]{1,0,0}6.5\pm 7.2}$ $-0.5\pm 1.3$ $5.7$
$C_{ud}^{(1)}$ $-2.4\pm 1.6$ $1.3$ ${\color[rgb]{1,0,0}-7.9\pm 7.2}$ ${\color[rgb]{1,0,0}-8.2\pm 6.8}$
@@
$C_{le}$ $-0.11\pm 0.31$ $3.5$ $-0.32\pm 0.41$ $2.9$ $-1.0\pm 3.6$ $3.5$ $-1.7\pm 3.6$ $3.3$
$C_{lu}$ $0.050\pm 0.084$ $6.4$ $6.5\pm 6.1$ $0.8$ $0.12\pm 0.85$ $7.4$ ${\color[rgb]{1,0,0}7.2\pm 7.9}$
$C_{ld}$ $-0.21\pm 0.24$ $3.7$ $2.1\pm 8.4$ $-1.3\pm 2.7$ $3.9$
$C_{qe}$ $0.057\pm 0.095$ $6.2$ $1.0\pm 3.2$ $1.1$ $0.0\pm 1.0$ $7.0$ ${\color[rgb]{1,0,0}6.6\pm 8.4}$
$C_{qu}^{(1)}$ $0.40\pm 0.80$ $2.1$ $6.8\pm 7.2$ $1.1\pm 6.3$ $2.7$ ${\color[rgb]{1,0,0}7.2\pm 7.9}$
$C_{qu}^{(8)}$ $-0.23\pm 0.55$ $2.7$ $-3.5\pm 2.7$ $1.0$ $-2.6\pm 5.4$ $2.7$ $-1.4\pm 7.0$ $2.7$
$C_{qd}^{(1)}$ $-3.5\pm 2.7$ $1.0$ ${\color[rgb]{1,0,0}-5.8\pm 9.3}$ ${\color[rgb]{1,0,0}5.5\pm 9.5}$
$C_{qd}^{(8)}$ $-5.0\pm 3.5$ $0.9$ $3.5\pm 9.1$ ${\color[rgb]{1,0,0}-6.8\pm 8.2}$
@@
$C$ $\Lambda/\sqrt{|C|}$ Obs $C$ $\Lambda/\sqrt{|C|}$ Obs $C$ $\Lambda/\sqrt{|C|}$ Obs
$C_{\phi\Box}$ $-0.7\pm 3.0$ $1.2$ H $-0.02\pm 0.29$ $1.3$ H
@@
$C_{ll}^{[aa33]}$ $-0.1\pm 3.7$ $3.7$ EW $0.00\pm 0.32$ $3.8$ EW $-0.002\pm 0.034$ $3.8$ EW
$C_{ll}^{[a33a]}$ $0.8\pm 3.2$ $3.7$ EW $0.06\pm 0.29$ $3.8$ EW $-0.003\pm 0.035$ $3.8$ EW
$C_{ll}^{[aabb]}$ $-2.5\pm 2.6$ $3.7$ EW $-0.21\pm 0.23$ $3.7$ EW $-0.026\pm 0.024$ $3.6$ EW
$C_{ll}^{[abba]}$ $-0.1\pm 1.0$ $6.8$ EW $-0.005\pm 0.099$ $6.7$ EW $-0.003\pm 0.011$ $6.3$ EW
$C_{qq}^{(1)[3333]}$ $-0.45\pm 0.23$ $10.7$ F $-0.043\pm 0.020$ $10.7$ F $-0.0036\pm 0.0018$ $11.9$ F
$C_{qq}^{(1)[aa33]}$ $0.211\pm 0.099$ $16.1$ F $0.0192\pm 0.0096$ $15.4$ F $0.00182\pm 0.00088$ $16.8$ F
$C_{qq}^{(1)[a33a]}$ $0.24\pm 0.13$ $14.4$ F $0.0188\pm 0.0096$ $15.3$ F $0.00178\pm 0.00086$ $17.0$ F
$C_{qq}^{(1)[aabb]}$ $-0.43\pm 0.22$ $10.2$ F $-0.036\pm 0.018$ $11.2$ F $-0.0035\pm 0.0016$ $12.2$ F
$C_{qq}^{(1)[abba]}$ $-0.47\pm 0.22$ $10.7$ F $-0.036\pm 0.021$ $10.1$ F $-0.0039\pm 0.0018$ $11.8$ F
$C_{qq}^{(3)[3333]}$ $-0.56\pm 0.23$ $10.2$ F $-0.040\pm 0.020$ $10.4$ F $-0.0038\pm 0.0018$ $11.9$ F
@@
$C_{qq}^{(3)[aabb]}$ $-0.36\pm 0.21$ $11.5$ F $-0.032\pm 0.019$ $11.2$ F $-0.0038\pm 0.0018$ $12.0$ F
$C_{qq}^{(3)[abba]}$ $-0.40\pm 0.21$ $10.6$ F $-0.036\pm 0.018$ $11.1$ F $-0.0039\pm 0.0019$ $11.3$ F
$C_{lq}^{(1)[3333]}$ ${\color[rgb]{1,0,0}-7.1\pm 7.1}$ EW $-0.9\pm 1.1$ $1.7$ EW $-0.21\pm 0.22$ $1.3$ F
$C_{lq}^{(1)[aa33]}$ $1.4\pm 1.6$ $4.6$ F $0.11\pm 0.15$ $4.6$ F $0.009\pm 0.017$ $4.7$ F
@@
$C_{ud}^{(1)[aa33]}$ $-5.1\pm 6.8$ DY
$C_{ud}^{(1)[33aa]}$ ${\color[rgb]{1,0,0}-7.4\pm 7.7}$ EW/T $-10.3\pm 3.7$ EW $-1.5\pm 1.9$ $0.4$ T
$C_{ud}^{(1)[aabb]}$ ${\color[rgb]{1,0,0}-6.0\pm 8.8}$ DY $-1.9\pm 2.6$ $1.1$ DY

### v2

@@ A.1 $U(3)^{5}$ flavour symmetric SMEFT
$C_{qe}^{[prst]}=C_{qe}\delta_{pr}\delta_{st}\,,\qquad C_{qu}^{(1),(8)[prst]}=C_{qu}^{(1),(8)}\delta_{pr}\delta_{st}\,,\qquad C_{qd}^{(1),(8)[prst]}=C_{qd}^{(1),(8)}\delta_{pr}\delta_{st}\,$,
$C_{quqd}^{(1),(8)[prst]}=C_{lequ}^{(1),(3)[prst]}=C_{ledq}^{[prst]}=0\,.$
@@ Appendix B Tables of results
In this appendix we collect a series of tables with the numerical results presented in the figures discussed in the main text, as well as some supplementary results. We report both the $68\%$ HPDI interval for the Wilson coefficients at the different scales considered in the study, as well as the bound on $\Lambda/\sqrt{|C_{i}|}$ in units of TeV obtained from the maximum of the $95\%$ HPDI interval for $|C_{i}|$. As indicated in the main text, in some cases the $68\%$ or $95\%$ HPDI (from which the bound on $\Lambda/\sqrt{|C_{i}|}$ is obtained) touches the edges of the prior dictated by perturbativity. These cases are indicated in the tables in red colour. On the other hand, in those cases where the posterior distribution of a given coefficient does not allow to set a bound, we indicate that using a dash.
@@ B.1 Results for the $U(3)^{5}$ flavour symmetric SMEFT
Table 7: Results of the individual fits for the $U(3)^{5}$ flavour symmetric SMEFT. For each coefficient, we report the $68\%$ HPDI interval, the bound on $\Lambda/\sqrt{|C_{i}|}$ in units of TeV obtained taking the maximum of the $95\%$ HPDI interval for $|C_{i}|$, and the set of observables that give the strongest constraint (EW stands for Electroweak, F for Flavour, H for Higgs, T for top, and DY for Drell-Yan). Results are shown with RG evolution for $\Lambda=10$ TeV, for $\Lambda=3$ TeV and without RG evolution. A dash denotes that no bound can be set from that particular fit. The meaning of the colour code is discussed in the text.
@@ 10 TeV 3 TeV 1 TeV - noRGE
$C_{i}$ $\Lambda/\sqrt{|C_{i}|}$ Obs $C_{i}$ $\Lambda/\sqrt{|C_{i}|}$ Obs $C_{i}$ $\Lambda/\sqrt{|C_{i}|}$ Obs
$C_{\phi\Box}$ $0.7\pm 3.0$ $1.2$ H $0.02\pm 0.29$ $1.3$ H
$C_{\phi D}$ $0.3\pm 1.1$ $6.4$ EW $0.040\pm 0.090$ $6.5$ EW $0.0024\pm 0.0072$ $7.7$ EW
$C_{\phi G}$ $0.112\pm 0.084$ $19.0$ H $0.0102\pm 0.0084$ $18.5$ H $0.0017\pm 0.0012$ $15.8$ H
$C_{\phi B}$ $0.10\pm 0.17$ $15.4$ H $0.008\pm 0.015$ $15.6$ H $0.0008\pm 0.0014$ $17.0$ H
$C_{\phi W}$ $0.33\pm 0.52$ $8.7$ H $0.028\pm 0.047$ $8.8$ H $0.0029\pm 0.0051$ $8.8$ H
$C_{G}$ $\color[rgb]{1,0,0}10.3\pm 4.7$} H/T $3.5\pm 1.8$ $1.1$ T $0.07\pm 0.15$ $1.7$ T
$C_{W}$ $2.2\pm 3.8$ $3.2$ H $0.25\pm 0.43$ $2.9$ H $0.50\pm 0.70$ $0.7$ EW
$C_{\phi l}^{(1)}$ $0.33\pm 0.45$ $9.1$ EW $0.030\pm 0.039$ $9.2$ EW $0.0029\pm 0.0039$ $9.8$ EW
$C_{\phi l}^{(3)}$ $0.67\pm 0.64$ $7.3$ EW $0.058\pm 0.056$ $7.3$ EW $0.0061\pm 0.0058$ $7.6$ EW
@@
$C_{\phi d}$ $\color[rgb]{1,0,0}-10.3\pm 4.4$} EW $0.83\pm 0.51$ $2.2$ EW $0.087\pm 0.051$ $2.3$ EW
$C_{ll}^{[aabb]}$ $1.7\pm 2.1$ $4.2$ EW $0.15\pm 0.19$ $4.1$ EW $0.019\pm 0.021$ $4.0$ EW
$C_{ll}^{[abba]}$ $0.01\pm 0.96$ $7.3$ EW $0.005\pm 0.091$ $6.9$ EW $0.004\pm 0.010$ $6.3$ EW
$C_{qq}^{(1)[aabb]}$ $5.8\pm 4.6$ $2.7$ EW $0.62\pm 0.54$ $2.3$ EW $1.00\pm 0.55$ $0.7$ T
@@
$C_{i}$ $\Lambda/\sqrt{|C_{i}|}$ $C_{i}$ $\Lambda/\sqrt{|C_{i}|}$ $C_{i}$ $\Lambda/\sqrt{|C_{i}|}$ $C_{i}$ $\Lambda/\sqrt{|C_{i}|}$
$C_{\phi\Box}$ $0.7\pm 3.0$ $1.2$ $4.6\pm 5.3$ $0.8$
$C_{\phi D}$ $0.040\pm 0.090$ $6.5$ $0.2\pm 4.4$ $1.0$ $0.3\pm 1.1$ $6.4$
$C_{\phi G}$ $0.0102\pm 0.0084$ $18.5$ $0.016\pm 0.013$ $14.1$ $0.112\pm 0.084$ $19.0$ $0.14\pm 0.10$ $16.8$
$C_{\phi B}$ $0.008\pm 0.015$ $15.6$ $0.17\pm 0.56$ $2.7$ $0.10\pm 0.17$ $15.4$ $0.5\pm 2.4$ $4.6$
$C_{\phi W}$ $0.028\pm 0.047$ $8.8$ $1.8\pm 1.5$ $1.4$ $0.33\pm 0.52$ $8.7$ $\color[rgb]{1,0,0}-12.0\pm 2.8$}
$C_{\phi WB}$ $0.018\pm 0.025$ $11.6$ $\pm 1.7 1.41.4$ $0.19\pm 0.29$ $11.3$ $7.3\pm 4.9$
$C_{G}$ $3.5\pm 1.8$ $1.1$ $2.4\pm 2.1$ $1.2$ $\color[rgb]{1,0,0}10.3\pm 4.7$} $\color[rgb]{1,0,0}8.2\pm 6.8$}
$C_{W}$ $0.25\pm 0.43$ $2.9$ $1.1\pm 7.0$ $0.8$ $2.2\pm 3.8$ $3.2$
$C_{\phi l}^{(1)}$ $0.030\pm 0.039$ $9.2$ $0.2\pm 1.5$ $1.7$ $0.33\pm 0.45$ $9.1$ $2.9\pm 3.3$ $3.5$
$C_{\phi l}^{(3)}$ $0.058\pm 0.056$ $7.3$ $0.80\pm 0.35$ $2.5$ $0.67\pm 0.64$ $7.3$ $4.8\pm 3.0$ $3.1$
$C_{\phi e}$ $0.052\pm 0.051$ $7.7$ $1.0\pm 2.5$ $1.3$ $0.62\pm 0.59$ $7.6$ $\color[rgb]{1,0,0}9.8\pm 5.1$}
$C_{\phi q}^{(1)}$ $0.10\pm 0.13$ $5.0$ $3.5\pm 1.3$ $1.2$ $0.8\pm 1.4$ $5.4$ $3.1\pm 4.7$ $2.9$
$C_{\phi q}^{(3)}$ $0.086\pm 0.052$ $6.8$ $0.61\pm 0.25$ $2.9$ $1.01\pm 0.58$ $6.7$ $6.4\pm 2.1$ $3.1$
$C_{\phi u}$ $0.13\pm 0.18$ $4.3$ $1.0\pm 2.3$ $1.3$ $1.3\pm 1.7$ $4.7$ $\color[rgb]{1,0,0}-6.8\pm 7.2$}
$C_{\phi d}$ $0.83\pm 0.51$ $2.2$ $3.9\pm 3.0$ $1.0$ $\color[rgb]{1,0,0}-10.3\pm 4.4$} $\color[rgb]{1,0,0}-8.2\pm 6.8$}
$C_{ll}^{[aabb]}$ $0.15\pm 0.19$ $4.1$ $1.7\pm 2.1$ $4.2$
$C_{ll}^{[abba]}$ $0.005\pm 0.091$ $6.9$ $0.82\pm 0.74$ $2.1$ $0.01\pm 0.96$ $7.3$ $2.1\pm 3.4$ $3.6$
$C_{qq}^{(1)[aabb]}$ $0.62\pm 0.54$ $2.3$ $11.2\pm 3.8$ $5.8\pm 4.6$ $2.7$ $\color[rgb]{1,0,0}-8.6\pm 6.5$}
$C_{qq}^{(1)[abba]}$ $0.09\pm 0.23$ $4.0$ $0.8\pm 1.0$ $1.8$ $0.3\pm 2.2$ $4.7$ $5.3\pm 3.4$ $2.9$
$C_{qq}^{(3)[aabb]}$ $0.034\pm 0.050$ $8.3$ $6\pm 0.33 3.03.0$ $0.21\pm 0.40$ $9.9$ $0.5\pm 3.1$ $3.8$
$C_{qq}^{(3)[abba]}$ $0.25\pm 0.16$ $4.0$ $0.24\pm 0.74$ $2.4$ $2.0\pm 1.4$ $4.5$ $7\pm 2.5 3.3$3.3
$C_{lq}^{(1)}$ $0.055\pm 0.054$ $7.4$ $2.5\pm 4.3$ $1.0$ $0.23\pm 0.51$ $8.8$ $3.7\pm 4.9$ $2.7$
$C_{lq}^{(3)}$ $0.0035\pm 0.0055$ $24.9$ $0.10\pm 0.14$ $4.9$ $0.028\pm 0.058$ $26.7$ $0.03\pm 0.51$ $9.7$
$C_{ee}$ $0.10\pm 0.12$ $5.3$ $4.6\pm 5.4$ $0.9$ $1.0\pm 1.3$ $5.3$ $2.9\pm 6.8$ $2.7$
$C_{uu}^{[aabb]}$ $0.47\pm 0.56$ $2.4$ $9.5\pm 5.6$ $2.2\pm 4.4$ $3.1$
$C_{uu}^{[abba]}$ $0.51\pm 0.40$ $2.6$ $2.3\pm 2.0$ $1.2$ $4.2\pm 4.1$ $2.8$ $\color[rgb]{1,0,0}-8.9\pm 6.1$}
@@
$C_{dd}^{[abba]}$ $3.9\pm 6.0$
$C_{eu}$ $0.036\pm 0.038$ $9.0$ $\pm 6.1 0.90.9$ $0.17\pm 0.41$ $10.2$ $2.6\pm 8.2.5$
$C_{ed}$ $0.10\pm 0.12$ $5.2$ $\color[rgb]{1,0,0}6.5\pm 7.2$} $0.5\pm 1.3$ $5.7$
$C_{ud}^{(1)}$ $2.4\pm 1.6$ $1.3$ $\color[rgb]{1,0,0}-7.9\pm 7.2$} $\color[rgb]{1,0,0}-8.2\pm 6.8$}
@@
$C_{le}$ $0.11\pm 0.31$ $3.5$ $0.28\pm 0.39$ $2.9$ $1.0\pm 3.6$ $3.5$ $1.7\pm 3.6$ $3.3$
$C_{lu}$ $0.050\pm 0.084$ $6.4$ $5.3\pm 6.7$ $0.8$ $0.12\pm 0.85$ $7.4$
$C_{ld}$ $0.21\pm 0.24$ $3.7$ $1.3\pm 2.7$ $3.9$
$C_{qe}$ $0.057\pm 0.095$ $6.2$ $1.6\pm 5.8$ $0.8$ $0.0\pm 1.0$ $7.0$
$C_{qu}^{(1)}$ $0.40\pm 0.80$ $2.1$ $6.8\pm 7.2$ $1.1\pm 6.3$ $2.7$
$C_{qu}^{(8)}$ $0.27\pm 0.53$ $2.6$ $3.7\pm 2.6$ $1.0$ $2.6\pm 5.4$ $2.7$ $1.8\pm 7.0$ $2.7$
$C_{qd}^{(1)}$ $0.1\pm 2.3$ $1.4$ $\color[rgb]{1,0,0}5.5\pm 9.5$}
$C_{qd}^{(8)}$ $4.8\pm 3.5$ $0.9$ $\color[rgb]{1,0,0}-6.8\pm 8.2$}
@@
$C_{i}$ $\Lambda/\sqrt{|C_{i}|}$ Obs $C_{i}$ $\Lambda/\sqrt{|C_{i}|}$ Obs $C_{i}$ $\Lambda/\sqrt{|C_{i}|}$ Obs
$C_{\phi\Box}$ $0.7\pm 3.0$ $1.2$ H $0.02\pm 0.29$ $1.3$ H
@@
$C_{ll}^{[aa33]}$ $0.1\pm 3.7$ $3.7$ EW $0.00\pm 0.32$ $3.8$ EW $0.002\pm 0.034$ $3.8$ EW
$C_{ll}^{[a33a]}$ $0.8\pm 3.2$ $3.7$ EW $0.06\pm 0.29$ $3.8$ EW $0.003\pm 0.035$ $3.8$ EW
$C_{ll}^{[aabb]}$ $2.5\pm 2.6$ $3.7$ EW $0.21\pm 0.23$ $3.7$ EW $0.026\pm 0.024$ $3.6$ EW
$C_{ll}^{[abba]}$ $0.1\pm 1.0$ $6.8$ EW $0.005\pm 0.099$ $6.7$ EW $0.003\pm 0.011$ $6.3$ EW
$C_{qq}^{(1)[3333]}$ $0.45\pm 0.23$ $10.7$ F $0.043\pm 0.020$ $10.7$ F $0.0036\pm 0.0018$ $11.9$ F
$C_{qq}^{(1)[aa33]}$ $0.235\pm 0.099$ $15.1$ F $0.0192\pm 0.0096$ $15.4$ F $0.00182\pm 0.00088$ $16.8$ F
$C_{qq}^{(1)[a33a]}$ $0.214\pm 0.096$ $15.5$ F $0.0188\pm 0.0096$ $15.3$ F $0.00178\pm 0.00086$ $17.0$ F
$C_{qq}^{(1)[aabb]}$ $0.43\pm 0.22$ $10.2$ F $0.036\pm 0.018$ $11.2$ F $0.0035\pm 0.0016$ $12.2$ F
$C_{qq}^{(1)[abba]}$ $0.47\pm 0.22$ $10.7$ F $0.036\pm 0.021$ $10.1$ F $0.0039\pm 0.0018$ $11.8$ F
$C_{qq}^{(3)[3333]}$ $0.56\pm 0.23$ $10.2$ F $0.040\pm 0.020$ $10.4$ F $0.0038\pm 0.0018$ $11.9$ F
@@
$C_{qq}^{(3)[aabb]}$ $0.36\pm 0.21$ $11.5$ F $0.032\pm 0.019$ $11.2$ F $0.0038\pm 0.0018$ $12.0$ F
$C_{qq}^{(3)[abba]}$ $0.40\pm 0.21$ $10.6$ F $0.036\pm 0.018$ $11.1$ F $0.0039\pm 0.0019$ $11.3$ F
$C_{lq}^{(1)[3333]}$ $\color[rgb]{1,0,0}-7.1\pm 7.1$} EW $0.9\pm 1.1$ $1.7$ EW $0.21\pm 0.24$ $1.2$ F
$C_{lq}^{(1)[aa33]}$ $0.21\pm 0.38$ $10.1$ F $0.019\pm 0.035$ $10.1$ F $0.0021\pm 0.0036$ $10.3$ F
@@
$C_{ud}^{(1)[aa33]}$ $5.1\pm 6.8$ DY
$C_{ud}^{(1)[33aa]}$ $\color[rgb]{1,0,0}-7.4\pm 7.7$} EW/T $10.3\pm 3.7$ EW $1.5\pm 1.9$ $0.4$ T
$C_{ud}^{(1)[aabb]}$ $\color[rgb]{1,0,0}-6.0\pm 8.8$} DY $9\pm 2.6 1.1$1.1 DY